OpenStudy (anonymous):

If f(x)=3+x/4+5/x^2, find f'(x).

OpenStudy (anonymous):

you can rewrite it as f(x) = 3 + x/4 +5x^(-2) and just use the power rule to get f'(x)

OpenStudy (aravindg):

Use the property that derivative of sum=sum of derivatives

OpenStudy (anonymous):

If f(x)=3+x/4+5/x^2 then f'(x)=3' + (x/4)' + [5/(x^2)]' => f'(x)= 0+(1/4)+(-10/x^3)

OpenStudy (anonymous):

okay thanks

OpenStudy (aravindg):

yw :)